If train A leaves the station going 60 miles per hour and train B leaves one hour later going 85 miles per hour, how long will it take train B to catch up with train A?
Answer by checkley77(12569)

D=RT
D=60T FOR TRAIN A.
D=85(T-1) FOR THRAIN B.
KNOWING THAT THE DISTANCES ARE THE SAME WE SET THE DISTANCE EQUATIONS EQUAL TO EACH OTHER.
60T=85(T-1)
60T=85T-85
60T-85T=-85
-25T=-85
T=-85/-25
T=3.4 HOURS AFTER TRAIN A LEAVES TRAIN B WILL CATCH UP TO IT.
PROOF:
60*3.4=85(3.4-1)
204=85*2.4
204=204